OpenStudy (anonymous):

find the partial fraction decomposition: (4x+23)/(x^2-x-6)

OpenStudy (amistre64):

youll have to factor the denominator

OpenStudy (amistre64):

\[\frac{(4x+23)}{(x^2-x-6)}=\frac{(4x+23)}{(x+a)(x+b)}=\frac{A}{(x+a)}+\frac{B}{(x+b)}\]

OpenStudy (amistre64):

then you solve for A and B as such: \[4x+23=A(x+b)+B(x+a)\] when x=-a we get \[4(-a)+23=A(-a+b)+B(-a+a)\] \[-4a+23=A(-a+b)\] \[\frac{-4a+23}{-a+b}=A\]

OpenStudy (amistre64):

same concept for solving for B; just make x=-b